### v1

@@ Abstract
| Gaia ID | RA | Dec | Star Name | G | Distance | BP - RP | Radial Velocity | MJD | Exposure | Source | Notes |
| (h:m:s) | (d:m:s) | (mag, | (kpc) | (mag, | (km s) | Time (s) | |||||
@@
The 300S stream was first found unintentionally by Geha et al. (2009) while studying the dwarf galaxy Segue 1, which overlaps 300S on the sky. Its stars have since been re-observed spectroscopically by the Southern Stellar Stream Spectroscopic Survey, S (Li et al., 2019, 2022). From its highly eccentric and retrograde orbit, it is clear the stream’s progenitor was accreted onto the Milky Way from another galaxy (Fu et al., 2018). 300S’s progenitor may have formed in the galaxy Gaia-Sausage-Enceladus (GSE, Belokurov et al. 2018; Helmi et al. 2018) prior to its merger with the Milky Way (Fu et al., 2018; Li et al., 2022).
@@ 2 Observations
We selected ten of the brightest candidate member stars of 300S in three ways (Table 1). First, we chose the six brightest 300S member stars found by S in Li et al. (2022). We refer readers to Li et al. (2022) for more details on the initial S data reduction and analysis, as well as the member star selection. As S limited the targets to for 300S stream to avoid cross-talk between the brightest and faintest targets, we further selected three possible stars using Gaia DR3 (Gaia Collaboration et al., 2016; Gaia Collaboration et al., 2022) at a brighter magnitude (or beyond the S footprint) which could be potentially missed by S, guided by the locations in the proper motion space and color-magnitude diagram of the previously identified 300S member stars from Li et al. (2022). Finally, we also included the most metal-poor star from Fu et al. (2018), as this star is the main driver of the metallicity dispersion revealed by Fu et al. (2018).
These ten stars were observed using the Magellan/MIKE high-resolution spectrograph on May 5th and 16th, 2021, using a 0.7 slit and 2 × 2 on-chip binning (Bernstein et al., 2003). We achieve a resolution of R 28,000 on the blue echelle orders and R 22,000 on the red echelle orders. The data were reduced using CarPy (Kelson, 2003). Details about the observations, including heliocentric radial velocities and exposure times, can be found in Table 1. The star J10201555 was also observed on April 13th, 2023 to increase the SNR, and the star J10121623 was observed on March 12th, 2022 to obtain an additional radial velocity. We achieve SNRs measured per pixel ranging from 60 to 100 for red wavelengths around 8,000 Å and 35 to 50 for blue wavelengths around 4,500 Å.
Fig. 1 depicts the position and velocities of stars in the same region of sky as 300S. We calculate the radial velocity by cross-correlating each echelle order with HD122563. We take a weighted average of the orders’ velocities to find the radial velocity of the star, down-weighting orders with high uncertainty. The standard deviation of these velocities are taken to be the overall uncertainty of the star’s radial velocity. We also include radial velocities of stars obtained with S(Li et al., 2019) and Gaia DR3 (Katz et al., 2023). Of the ten stars observed with MIKE, two were found to have radial velocities inconsistent with the stream track, indicating they are not members of 300S and were therefore rejected from this study. In particular, there are four radial velocities for the most metal-poor candidate member star analyzed by Fu et al. (2018) (J10121623), showing clear velocity variations consistent with being in a close binary. We observed this star twice with MIKE to confirm its binarity. Though more velocities would be needed to robustly obtain a systemic velocity, the current data suggest that the SEGUE velocity was obtained at an unlucky phase of the binary orbit. Thus, we now classify this star as a radial velocity non-member.
@@ 3 Stellar Parameter Determination
Here, C is the Gaia color (G-RP) of the star. We use [Fe/H] as measured spectroscopically for each star as given in Table 2. We assume the star is in local thermodynamic equilibrium (LTE) and do not calculate non-LTE corrections (nLTE). Most of our stars are in the range of 4000 K to 4800 K. The most luminous star, J10451408, is slightly below this range at 3880 K.
@@ 4 Abundance Analysis
The measured abundances described here are shown in Table 3.
@@ 4.1 Abundance Dispersion Analysis
We calculate the intrinsic dispersion of all measured elements in the stars analyzed here. We calculate the mean abundance and dispersion of the stream given our measurements, using Eqn. 4. and represent the intrinsic dispersion and the individual error measurements, respectively, while is the mean abundance value. The log likelihood is:
@@ 5 Stream Progenitor Mass Estimate
We estimated the stellar mass of the 300S stream by comparing the number of confirmed red giant branch member stars to that of theoretical isochrones and their luminosity functions. We did this for all eight stars analyzed in this work, as well as the remaining 34 300S member stars from S(Li et al., 2022). The S observations are sufficiently complete within its observation footprint to a Gaia G magnitude of 17.2, as confirmed by matching the stream’s observations with the theoretical luminosity function.
We used the Dartmouth Stellar Evolution Database (Dotter et al., 2008)555http://stellar.dartmouth.edu/models/ to generate isochrones and visually fit them to our data. The best-fit isochrone and the isochrone using stream parameters as measured by Fu et al. (2018) and Li et al. (2022) are shown in Fig. 2. We fit DECam DECaLs DR9 photometry for all 300S member stars observed by S(Dey et al., 2019). We created a background-subtracted Hess diagram to best match the isochrones to the main sequence turn-off. From the isochrones, we find the cluster to be at a distance kpc, have an age of 12.5 1.5 Gyr, [Fe/H] = , and [/Fe] = 0.0.
@@
From this calculation, we find that the current mass of 300S is 10 . However, we did not include stars which extend beyond S’s footprint for the stream. 10 is therefore a lower limit for the progenitor mass. More sophisticated methods of calculating masses could be used in future investigations (e.g., Martin et al., 2008, 2016).
@@
Here, , , and are the distance at apocenter in kpc, the circular velocity of the Milky Way (assumed to be 240 km s) and the eccentricity of the system’s orbit, respectively. Since 300S is completely dissolved and we can find no intact cluster, we use known orbital parameters and possible dissolution times to see the highest and lowest possible initial masses that would be disrupted by now. We use the cluster’s lower age limit and the age of the universe as minimum and maximum dissolution times respectively. The cluster can disrupt any time after forming, so this is essentially a range for the upper limit of the dissolution time. We used 300S’s orbital parameters found in Li et al. (2022) and the best-fit age we found using isochrones. We find the maximum initial mass of 300S to be M = 10 .
In conclusion, we find the progenitor mass of 300S is likely to be in the range of 10 . We note, however, that models of mass loss are under development and constraints on the initial mass may improve significantly with time.
@@ 6 Results
We analyzed ten red giant branch member candidates of the 300S stream. We found two of the stars to be inconsistent with stream members (including the most metal-poor star analyzed in Fu et al. 2018), due to discrepant radial velocities. The eight analyzed stars had a mean LTE metallicity of [Fe/H] = with an intrinsic dispersion of less than 0.09 at 95% confidence level. Abundances measured for each star can be found in Table 3. Mean element abundances and dispersions are listed in Table 4. Of the eight stars, we highlight three in particular: the enriched J10201555, the CH star J10011554, and the cool star J10451408.
@@ 7.1 Globular Cluster, Not a Dwarf Galaxy
Previous studies of 300S struggled to determine whether the progenitor of the stellar stream was a dwarf galaxy or a globular cluster. The stream was found by Fu et al. (2018) to have a metallicity dispersion. Li et al. (2022), however, measured no metallicity dispersion. To settle the contradicting results, we re-observed the most metal-poor star in Fu et al. (2018) with MIKE twice to verify its membership. The star’s measured radial velocity is inconsistent with the stream moving at 300 km s; further, the star shows significant variations in its radial velocities over time: on the two occasions the star was observed with MIKE, the observed radial velocity was measured to be km s. Smeasured a velocity of km s. Lastly, the SEGUE data used by Fu et al. (2018) indicates a velocity of km s. These large variations in radial velocity (>50 km s) measured over several years imply the star is a part of a binary system. The average velocity is not similar enough to the stellar stream to be considered a member, as can be observed in Figure 1. Excluding this star from the analysis, the 300S stars have an insignificant metallicity spread.
@@ 7.2 Multiple Populations and Initial Mass
We further confirm that the progenitor of 300S was a globular cluster by detecting a second-population star, J10201555. However, the stream does not contain as many enriched stars as we had initially expected. A correlation exists between the mass of intact Milky Way GCs and the fraction of enriched stars found therein, as seen in the left panel of Fig. 8. Assuming a binomial distribution, we infer that for just one of eight stars to be enriched, the stream is likely to have an inherent enrichment fraction of f, as is reflected in the figure. If we choose to exclude the cool star entirely from the sample due to its ambiguous enrichment, we calculate the stream’s inherent enrichment fraction to be f. We estimated 300S’s current stream mass to be 10, as described in Section 5. The total stream mass should be higher than this, since more stream members beyond the footprint of S were not included in this estimate. Interpolating a linear relationship between the current mass of intact Milky Way GCs and their fraction of enriched stars (as seen in the left panel of Fig. 8), we expected half of the stars in our sample to be enriched. Given our small sample size, we cannot completely exclude a 50% enrichment fraction, but we proceed taking the observed enriched fraction at face value.
@@
With just one of the eight stars being enriched, 300S appears underenriched when compared to the present day masses of GCs, but is fairly reasonable when we compare the progenitor mass to the initial masses of intact Milky Way GCs. Intact Milky Way GCs have lost, on average, 80% of their initial mass over their lifetimes (Baumgardt et al., 2019) due to a combination of evaporation and tidal stripping. In Section 5 we estimated limits on the progenitor mass of 300S to be in the range of 10. In the right panel of Fig. 8, we compare the initial masses of intact Milky Way GCs with our best constraints on the initial mass of 300S. 300S does not appear anomalous in this framework, as it appears relatively in line with the relationship of enrichment fraction to initial mass as can be seen in other clusters containing MSPs (Belokurov & Kravtsov, 2023). The stream progenitor could still be considered underenriched if we assume a linear relationship between fractional enrichment and initial mass of the globular cluster, but it could also indicate a non-linear relationship between the cluster parameters.
@@
The masses of GCs in the Large and Small Magellanic Clouds (represented by black stars in Fig. 8, as measured by Mucciarelli et al. 2009; Mateluna et al. 2012; Mucciarelli et al. 2014; Hollyhead et al. 2017; Niederhofer et al. 2017a, b; Martocchia et al. 2017; Hollyhead et al. 2018; Zhang et al. 2018; Martocchia et al. 2018, and Hollyhead et al. 2019 and compiled by Gratton et al. 2019) support this theory. It has been suggested that these clusters may have lost less mass, since the Magellanic Clouds are significantly less massive than the Milky Way (e.g., Milone et al., 2020). We therefore assume that the Magellanic Cloud GC initial masses are relatively similar to the masses we see now. These clusters’ masses fall in line with calculated initial masses of intact Milky Way GCs, and trend towards having a non-linear increase in enrichment above 10.
@@
It has been suggested a minimum critical mass must be achieved in order to create MSPs (e.g., Gratton et al., 2019), though this threshold mass is currently undetermined. The initial mass of 300S may hover near the critical threshold, allowing the unusually low enrichment we see here. Further improvements to modelling of GCs to better calculate initial masses will allow us to more strongly constrain the mass-enrichment relationship. Additionally, spectroscopic follow-up of more globular cluster stellar streams may further serve as a useful probe to finding a critical mass needed to achieve chemical enrichment found in most GCs. Theoretically, Bland-Hawthorn et al. (2010) suggests that a single-population globular cluster is achievable up to an initial mass of 10 . If this corresponds to the minimum mass threshold, it would suggest significant revisions are needed to globular cluster mass loss.
@@ 8 Conclusion
The progenitor of 300S straddled the threshold of the critical mass for making MSPs. We calculate the initial mass of 300S based on its orbital parameters and compare it to the initial masses of intact GCs as calculated by Baumgardt et al. (2019). We find it to have an initial mass of M = 10 , which aligns with the initial masses of GCs with small fractions of enriched stars. This comparison is shown in Fig. 8. 300S is therefore a good benchmark for the critical mass threshold of MSPs.

### v2

@@ Abstract
| Gaia ID | RA | Dec | Star Name | G0 | Distance | BP - RP | Radial Velocity | MJD | Exposure | Source | Notes |
| (∘) | (∘) | (mag, | (kpc) | (mag, | (km s-1) | Time (s) | |||||
@@
The 300S stream was first found unintentionally by Geha et al. (2009) while studying the dwarf galaxy Segue 1, which overlaps 300S on the sky. Its stars have since been re-observed spectroscopically by the Southern Stellar Stream Spectroscopic Survey, S5 (Li et al., 2019, 2022). From its highly eccentric and retrograde orbit, it is clear the stream’s progenitor was accreted onto the Milky Way from another galaxy (Fu et al., 2018). 300S’s progenitor may have formed in the galaxy Gaia-Sausage-Enceladus (GSE, Belokurov et al. 2018; Helmi et al. 2018) prior to its merger with the Milky Way (Fu et al., 2018; Li et al., 2022).
@@ 2 Observations
We selected ten of the brightest candidate member stars of 300S in three ways (Table 1). First, we chose the six brightest 300S member stars found by S5 in Li et al. (2022). We refer readers to Li et al. (2022) for more details on the initial S5 data reduction and analysis, as well as the member star selection. As S5 limited the targets to for 300S stream to avoid cross-talk between the brightest and faintest targets, we further selected three possible stars using Gaia DR3 (Gaia Collaboration et al., 2016; Gaia Collaboration et al., 2022) at a brighter magnitude (or beyond the S5 footprint) which could be potentially missed by S5, guided by the locations in the proper motion space and color-magnitude diagram of the previously identified 300S member stars from Li et al. (2022). Finally, we also included the most metal-poor star from Fu et al. (2018), as this star is the main driver of the metallicity dispersion revealed by Fu et al. (2018).
These ten stars were observed using the Magellan/MIKE high-resolution spectrograph on May 5th and 16th, 2021, using a 0.7′′ slit and 2 × 2 on-chip binning (Bernstein et al., 2003). We achieve a resolution of R 28,000 on the blue echelle orders and R 22,000 on the red echelle orders. The data were reduced using CarPy (Kelson, 2003). Details about the observations, including heliocentric radial velocities and exposure times, can be found in Table 1. The star J10201555 was also observed on April 13th, 2023 to increase the SNR, and the star J10121623 was observed on March 12th, 2022 to obtain an additional radial velocity. We achieve SNRs measured per pixel ranging from 60 to 100 for red wavelengths around 8,000 Å and 35 to 50 for blue wavelengths around 4,500 Å.
Fig. 1 depicts the position and velocities of stars in the same region of sky as 300S. We calculate the radial velocity by cross-correlating each echelle order with HD122563. We take a weighted average of the orders’ velocities to find the radial velocity of the star, down-weighting orders with high uncertainty. The standard deviation of these velocities are taken to be the overall uncertainty of the star’s radial velocity. We also include radial velocities of stars obtained with S5(Li et al., 2019) and Gaia DR3 (Katz et al., 2023). Of the ten stars observed with MIKE, two were found to have radial velocities inconsistent with the stream track, indicating they are not members of 300S and were therefore rejected from this study. In particular, there are four radial velocities for the most metal-poor candidate member star analyzed by Fu et al. (2018) (J10121623), showing clear velocity variations consistent with being in a close binary. We observed this star twice with MIKE to confirm its binarity. Though more velocities would be needed to robustly obtain a systemic velocity, the current data suggest that the SEGUE velocity was obtained at an unlucky phase of the binary orbit. Thus, we now classify this star as a radial velocity non-member.
@@ 3 Stellar Parameter Determination
Here, C is the Gaia color (G-RP)0 of the star. We use [Fe/H] as measured spectroscopically for each star as given in Table 2. We assume the star is in local thermodynamic equilibrium (LTE) and do not calculate non-LTE corrections (nLTE). Most of our stars are in the range of 4000 K to 4800 K. The most luminous star, J10451408, is slightly below this range at 3880 K.
@@ 4 Abundance Analysis
The measured abundances described here are shown in Table LABEL:tab:abundsummary.
@@ 4.1 Abundance Dispersion Analysis
We calculate the intrinsic dispersion of all measured elements in the stars analyzed here. We calculate the mean abundance and dispersion of the stream given our measurements, using Eqn. 3. and represent the intrinsic dispersion and the individual error measurements, respectively, while is the mean abundance value. The log likelihood is:
@@ 5 Stream Progenitor Mass Estimate
We estimated the stellar mass of the 300S stream by comparing the number of confirmed red giant branch member stars to that of theoretical isochrones and their luminosity functions. We did this for all eight stars analyzed in this work, as well as the remaining 34 300S member stars from S5(Li et al., 2022). The S5 observations are sufficiently complete within its observation footprint to a Gaia G magnitude of 17.2, as confirmed by matching the stream’s observations with the theoretical luminosity function.
We used the Dartmouth Stellar Evolution Database (Dotter et al., 2008)555http://stellar.dartmouth.edu/models/ to generate isochrones and visually fit them to our data. The best-fit isochrone and the isochrone using stream parameters as measured by Fu et al. (2018) and Li et al. (2022) are shown in Fig. 2. We fit DECam DECaLs DR9 photometry for all 300S member stars observed by S5(Dey et al., 2019). We created a background-subtracted Hess diagram to best match the isochrones to the main sequence turn-off. From the isochrones, we find the cluster to be at a distance kpc, have an age of 12.5 1.5 Gyr, [Fe/H] = , and [/Fe] = 0.0.
@@
From this calculation, we find that the current mass of 300S is 10 . However, we did not include stars which extend beyond S5’s footprint for the stream. 104.6 is therefore a lower limit for the progenitor mass. More sophisticated methods of calculating masses could be used in future investigations (e.g., Martin et al., 2008, 2016).
@@
Here, , , and are the distance at apocenter in kpc, the circular velocity of the Milky Way (assumed to be 240 km s-1) and the eccentricity of the system’s orbit, respectively. Since 300S is completely dissolved and we can find no intact cluster, we use known orbital parameters and possible dissolution times to see the highest and lowest possible initial masses that would be disrupted by now. We use the cluster’s lower age limit and the age of the universe as minimum and maximum dissolution times respectively. The cluster can disrupt any time after forming, so this is essentially a range for the upper limit of the dissolution time. We used 300S’s orbital parameters found in Li et al. (2022) and the best-fit age we found using isochrones. We find the maximum initial mass of 300S to be M = 104.8±0.1 .
In conclusion, we find the progenitor mass of 300S is likely to be in the range of 104.5-4.9 . We note, however, that models of mass loss are under development and constraints on the initial mass may improve significantly with time.
@@ 6 Results
We analyzed ten red giant branch member candidates of the 300S stream. We found two of the stars to be inconsistent with stream members (including the most metal-poor star analyzed in Fu et al. 2018), due to discrepant radial velocities. The eight analyzed stars had a mean LTE metallicity of [Fe/H] = with an intrinsic dispersion of less than 0.09 at 95% confidence level. Abundances measured for each star can be found in Table LABEL:tab:abundsummary. Mean element abundances and dispersions are listed in Table 4. Of the eight stars, we highlight three in particular: the enriched J10201555, the CH star J10011554, and the cool star J10451408.
@@ 7.1 Globular Cluster, Not a Dwarf Galaxy
Previous studies of 300S struggled to determine whether the progenitor of the stellar stream was a dwarf galaxy or a globular cluster. The stream was found by Fu et al. (2018) to have a metallicity dispersion. Li et al. (2022), however, measured no metallicity dispersion. To settle the contradicting results, we re-observed the most metal-poor star in Fu et al. (2018) with MIKE twice to verify its membership. The star’s measured radial velocity is inconsistent with the stream moving at 300 km s-1; further, the star shows significant variations in its radial velocities over time: on the two occasions the star was observed with MIKE, the observed radial velocity was measured to be km s-1. S5measured a velocity of km s-1. Lastly, the SEGUE data used by Fu et al. (2018) indicates a velocity of km s-1. These large variations in radial velocity (>50 km s-1) measured over several years imply the star is a part of a binary system. The average velocity is not similar enough to the stellar stream to be considered a member, as can be observed in Figure 1. Excluding this star from the analysis, the 300S stars have an insignificant metallicity spread.
@@ 7.2 Multiple Populations and Initial Mass
We further confirm that the progenitor of 300S was a globular cluster by detecting a second-population star, J10201555. However, the stream does not contain as many enriched stars as we had initially expected. A correlation exists between the mass of intact Milky Way GCs and the fraction of enriched stars found therein, as seen in the left panel of Fig. 8. Assuming a binomial distribution, we infer that for just one of eight stars to be enriched, the stream is likely to have an inherent enrichment fraction of f, as is reflected in the figure. If we choose to exclude the cool star entirely from the sample due to its ambiguous enrichment, we calculate the stream’s inherent enrichment fraction to be f. We estimated 300S’s current stream mass to be 104.5, as described in Section 5. The total stream mass should be higher than this, since more stream members beyond the footprint of S5 were not included in this estimate. Interpolating a linear relationship between the current mass of intact Milky Way GCs and their fraction of enriched stars (as seen in the left panel of Fig. 8), we expected half of the stars in our sample to be enriched. Given our small sample size, we cannot completely exclude a 50% enrichment fraction, but we proceed taking the observed enriched fraction at face value.
@@
With just one of the eight stars being enriched, 300S appears underenriched when compared to the present day masses of GCs, but is fairly reasonable when we compare the progenitor mass to the initial masses of intact Milky Way GCs. Intact Milky Way GCs have lost, on average, 80% of their initial mass over their lifetimes (Baumgardt et al., 2019) due to a combination of evaporation and tidal stripping. In Section 5 we estimated limits on the progenitor mass of 300S to be in the range of 104.5-4.9. In the right panel of Fig. 8, we compare the initial masses of intact Milky Way GCs with our best constraints on the initial mass of 300S. 300S does not appear anomalous in this framework, as it appears relatively in line with the relationship of enrichment fraction to initial mass as can be seen in other clusters containing MSPs (Belokurov & Kravtsov, 2023). The stream progenitor could still be considered underenriched if we assume a linear relationship between fractional enrichment and initial mass of the globular cluster, but it could also indicate a non-linear relationship between the cluster parameters.
@@
The masses of GCs in the Large and Small Magellanic Clouds (represented by black stars in Fig. 8, as measured by Mucciarelli et al. 2009; Mateluna et al. 2012; Mucciarelli et al. 2014; Hollyhead et al. 2017; Niederhofer et al. 2017a, b; Martocchia et al. 2017; Hollyhead et al. 2018; Zhang et al. 2018; Martocchia et al. 2018, and Hollyhead et al. 2019 and compiled by Gratton et al. 2019) support this theory. It has been suggested that these clusters may have lost less mass, since the Magellanic Clouds are significantly less massive than the Milky Way (e.g., Milone et al., 2020). We therefore assume that the Magellanic Cloud GC initial masses are relatively similar to the masses we see now. These clusters’ masses fall in line with calculated initial masses of intact Milky Way GCs, and trend towards having a non-linear increase in enrichment above 105.
@@
It has been suggested a minimum critical mass must be achieved in order to create MSPs (e.g., Gratton et al., 2019), though this threshold mass is currently undetermined. The initial mass of 300S may hover near the critical threshold, allowing the unusually low enrichment we see here. Further improvements to modelling of GCs to better calculate initial masses will allow us to more strongly constrain the mass-enrichment relationship. Additionally, spectroscopic follow-up of more globular cluster stellar streams may further serve as a useful probe to finding a critical mass needed to achieve chemical enrichment found in most GCs. Theoretically, Bland-Hawthorn et al. (2010) suggests that a single-population globular cluster is achievable up to an initial mass of 106.5 . If this corresponds to the minimum mass threshold, it would suggest significant revisions are needed to globular cluster mass loss.
@@ 8 Conclusion
The progenitor of 300S straddled the threshold of the critical mass for making MSPs. We calculate the initial mass of 300S based on its orbital parameters and compare it to the initial masses of intact GCs as calculated by Baumgardt et al. (2019). We find it to have an initial mass of M = 104.5-4.9 , which aligns with the initial masses of GCs with small fractions of enriched stars. This comparison is shown in Fig. 8. 300S is therefore a good benchmark for the critical mass threshold of MSPs.